Question 2

Steve Grant, the new controller of Bonita Ltd., has reviewed the expected useful lives and residual values of selected depreciable assets at the beginning of 2017. His findings are as follows.

Date

Accumulated
Depreciation

Useful life
in Years

Residual Value

Type of Asset

Acquired

Cost

1/1/17

Old

Proposed

Old

Proposed

Building 1/1/07 £810,900 £191,500 40 50 £44,900 £45,400
Warehouse 1/1/12 106,000 19,340 25 20 9,300 17,660


All assets are depreciated by the straight-line method. Bonita uses a calendar year in preparing annual financial statements. After discussion, management has agreed to accept Grant’s proposed changes.

Compute the revised annual depreciation on each asset in 2017.

Building

Warehouse

Revised annual depreciation £

£

Prepare the entry (or entries) to record depreciation on the building in 2017. (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Dec. 31
Answer #1
Computation of Revised Depreciaton
Revised Deprecation for Building = (810900-45400)/(50-10)= 19137.50
Revised Depreciation on Warehouse = (106000-17660)/(20-5)= $5889
Date Account Tittle Debit Credit
31-Dec Depreciaton on Building $19,137.50
Accumualted Depreciaiton- Buliding $19,137.50
31-Dec Depreciaton on Warehouse $5,889.00
Accumualted Depreciaiton- Warehouse $5,889.00
Answer #2

Revised Depreciation for Building

Cost

810900
Less: Accumulated Depreciation
-191500
Book Value
619400
Less: Salvage Value (Proposed)
-45400
Depreciable Cost
574000
Proposed Useful Life
50
Used Life
(2017-2007)=10

Revised Annual Depreciation

14350

Revised Annual Depreciation = 574000/40 = 14350


Revised Depreciation for Warehouse

Cost

106000
Less: Accumulated Depreciation
-19340
Book Value

86660

Less: Salvage Value (Proposed)
17660
Depreciable Cost
69000
Proposed Useful Life
20
Used Life
-5

Revised Annual Depreciation

15

Revised Annual Depreciation = 69000/15 = 4600



answered by: Basquemath
Answer #3

Steve Grant, the new controller of Bramble Ltd., has reviewed the expected useful lives and residual values of selected depreciable assets at the beginning of 2020. His findings are as follows.



Date




Accumulated
Depreciation


Total Useful life
in Years


Residual Value

Type of Asset


Acquired


Cost


1/1/20


Old


Proposed


Old


Proposed

Building
1/1/10
£804,000
£191,000
40
50
£40,000
£53,000
Warehouse
1/1/15
114,000
21,880
25
20
4,600
29,120